OpenStudy (anonymous):

Evaluate: 32 + (8 − 2) • 4 − 6/3

OpenStudy (anonymous):

PEMDAS order of operations :P

OpenStudy (hba):

Yeah Use BODMAS RULE

OpenStudy (hba):

OR PEMDAS

OpenStudy (hba):

Well BODMAS STANDS FOR BRACKETS OF DIVISION,MULTIPLICATION,ADDITION AND SUBRACTION

OpenStudy (hba):

So You Go According To The Order-Got It @rachelraver

OpenStudy (anonymous):

I can do it all until it gets into the fractions.

OpenStudy (hba):

Show Me Your Working

OpenStudy (hba):

@rachelraver

OpenStudy (goten77):

@gaara438125 6/3 is actually 2

OpenStudy (anonymous):

stop being smart!

OpenStudy (hba):

Yeah Right LoL Haha

OpenStudy (hba):

So @rachelraver Is Gonna Show Us How She Solves IT And Where She gets Stuck :)

OpenStudy (anonymous):

8 - 2 = 6 6 * 4 = 24 so 32 + 24 = 56 and since 6/3 = 2 the answer would be 54 but that's not an answer choice.

OpenStudy (hba):

Well Then What Is Your Answer Choice,Btw This Is The Right Answer :)

OpenStudy (anonymous):

nevermind, I typed it wrong! the 32 is supposed to be 3 to the second power. omg guys.

OpenStudy (hba):

the correct write the eq is 32+(8-2)*((4-6)/3) then 32+(6)*(-2/3)=32+(-12/3) =32-4=28

OpenStudy (anonymous):

so 9 + 22 = your answer

OpenStudy (anonymous):

- 2 28! finally.

OpenStudy (anonymous):

thank you guys.

OpenStudy (hba):

You Are Welcome @rachelraver :)

OpenStudy (anonymous):

32+(8-2)*4-6/3 Subtract 8 and 2 32+6*4-6/3 Multiply 6 and 4 32+24-6/3 divide 6 and 3 32+24-2 Add 32 and 24 56-2 Then just subtract Your answer should be 54 (if i am correct)

OpenStudy (anonymous):

she's had the answer for a while now, and she messed up the question

OpenStudy (anonymous):

ohhh opps! sorry.

OpenStudy (hba):

Haha LoL And Then I Worked It Out,Thanks To Me

OpenStudy (anonymous):

ok, but not really x]